OpenStudy (anonymous):

Help Please! Medals for help! Create a quadratic function, f(x), in vertex form. The a should be between 4 and –4, the h will be your birth month, and the k will be your birth day. Write your equation below. Using complete sentences, explain how to convert your birthday function into standard form. Graph your function. Include your graph below. Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7.

OpenStudy (anonymous):

My birthday is 10/7

OpenStudy (anonymous):

@campbell_st @Donblue22 @abb0t @katragaddasaichandra @StudyGurl14 @wesley747 @XxAminaxX @HelpBlahBlahBlah @vita955 @nerdgirl6 @lizzyrockefeller @lordofpens @thekakashigaming @Frosty_Flow @bahrom7893 @brucebaner @iheartfood @Jesstho.-. help?

OpenStudy (anonymous):

whos this?

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

how do you know my @

OpenStudy (anonymous):

i just type in @ and a letter then names of people online shows up

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

can you help me? @thekakashigaming

OpenStudy (anonymous):

.......................

OpenStudy (anonymous):

What up?

OpenStudy (anonymous):

Could someone help me with my question?

OpenStudy (campbell_st):

ok.... so of you are born on 30th june so you need a value of a.... you can choose.... the standard form I use is \[(x - h)^2 = 4a(y - k)\] wheere (h, k) is the vertex and a is the focal length.... using the the example date h = 6 and k = 30 then what you probably have to do is make y the subject. hope that makes sense

OpenStudy (anonymous):

:/ it helps a litte

OpenStudy (anonymous):

Could someone help me step by step?

OpenStudy (campbell_st):

you could be really cheeky and chose a = 0 ok... what is the birth month and birth day

OpenStudy (anonymous):

10/7

OpenStudy (campbell_st):

ok.... so based on the information the vertex is (birth month, Birth day) so ( 10, 7) which means h = 10 and k = 7 so using the standard form I posted above you have \[(x - 10)^2 = 4a(y - 7)\] next you need a value of a

OpenStudy (anonymous):

A would be between 4 through -4

OpenStudy (campbell_st):

so what are you choosing....?

OpenStudy (anonymous):

i'm going with 3

OpenStudy (campbell_st):

ok... so then you have \[(x - 10)^2= 12(y - 7)\] so now make y the subject if your vertex from is \[y = p(x - h)^2 + k\]

OpenStudy (anonymous):

do i use 4 for x

OpenStudy (campbell_st):

well using the form I posted above... divide both sides of the equation by 12 and then add 7... hope it makes sense

OpenStudy (anonymous):

:/ why divide by 12

OpenStudy (anonymous):

in my class, we use the formula f(x) = ax^2 + bx^2 + c

OpenStudy (anonymous):

but i dont understand it

OpenStudy (anonymous):

@campbell_st ?

OpenStudy (anonymous):

....

OpenStudy (anonymous):

anyone?

OpenStudy (campbell_st):

you want y as the subject so \[y - 7 = \frac{1}{12}(x - 10)^2\] so \[y = \frac{1}{12}(x - 10)^2 + 7\] which is a vertex form... expand the quadratic

OpenStudy (anonymous):

do you solve both sides of the equation? sorry if im asking too many questions

OpenStudy (anonymous):

can you help explain also, @Jesstho.-. and @Firejay5 ?

OpenStudy (anonymous):

......

OpenStudy (anonymous):

(-)_(-)

OpenStudy (nerdgirl6):

Hello @adoomyG maybe I can help you.

OpenStudy (nerdgirl6):

First, you know vertex form is y=a(x-h)^2+k So plug your birthday numbers into that.

OpenStudy (campbell_st):

you then expand the perfect square \[y = \frac{1}{12}(x^2 - 20x + 100) + 7\] now simplify this.